# Characterizing Density and Gravitational Potential Fluctuations of the Interstellar Medium

### v1

@@ 3.2 Two-Point Spatial Surface Density Statistics
Although we have presented Figure 6 for just the $\delta$ field in the R8 simulation, we note that analogous plots of $P_{s}(\mathbf{k})$ for the R8 simulation, and for both $P_{\delta}(\mathbf{k})$ and $P_{s}(\mathbf{k})$ for the LGR4 simulations, look qualitatively the same. In particular, they all show a near-isotropic field with a slight preference for trailing waves. Even for the scales that deviate most strongly from isotropy, the variation in the power at a fixed $k$ but varied $(k_{x},k_{y})$ is well within $\sim 0.25\,$dex, and so for the remainder of this paper, for simplicity, we will approximate all fluctuation power spectra to be a function of the magnitude of the wavevector, $k$, alone (but see Section 5.3 for further discussion).
Figure 7: Spatial power spectra of the linear fluctuations (left) and logarithmic fluctuations (right) for the R8 (purple) and LGR4 (orange) simulations. The solid curves of each color are the time-averaged spectra averaged over the azimuthal angle in $k$-space, while shaded regions show ±$\pm 1\sigma$ variations around each curve due to anisotropy (see, e.g., Figure 6) and time-dependence. The left and right shaded grey regions in each panel are excluded from fitting the power-law slope (indicated by the colored dashed curves) due to box-size and resolution effects respectively.
In Figure 7 we plot the power spectra in $\delta$ (left) and $s$ (right) for the R8 (purple) and LGR4 (orange) simulations as a function of $k$, calculated by averaging the spectra $P_{\delta}(\mathbf{k})$ and $P_{s}(\mathbf{k})$ over contours of constant $k$ (e.g., the white dashed circles in Figure 6). The solid curves all indicate the measured power spectra, while the shaded regions around them represent ±$\pm 1\sigma$ deviations in the values due to time-dependence in the fluctuation Fourier components and anisotropy. We see that the shaded regions are wider at smaller $k$, demonstrating the increased anisotropy present there, but the fractional variation remains small, illustrating that the approximation of isotropy is reasonable.
For both $P_{\delta}$ and $P_{s}$, the power spectra in both simulations are well-fit by power laws across a wide range of scales. We measure power law indices by fitting the slope of the curves over the range $kL\in[10,200]$; for smaller $kL$ box-size effects and anisotropy may become important, while for larger $kL$ there is contamination due to the proximity to the grid scale $\Delta x$. We demonstrate in Appendix A that our results for the largest scales, which carry the most power, are not affected by the simulation spatial resolution. Thus, our model for the power spectra takes the form
@@
We see from Figure 7 and Table 2 that despite the stark differences in the galactic environments and feedback strengths between the R8 and LGR4 simulations, their dimensionless fluctuation power spectra as functions of $kL$ have power law indices consistent with each other and even share similar normalizations. Because the power law spectral indices are sufficiently steep ($n_{\delta},n_{s}>2$), the power in these fluctuations is dominated by the largest scales. We speculate on the origin of the similarity across environments, as well as the broader implications of the steep power-law structure of the field in Section 5.1.
3.3 Two-Point Spatio-temporal Surface Density Statistics
So far, we have focused on the spatial features of the fluctuations, and averaged the relevant statistics over time as a proxy for an ensemble average. However, we are also able to probe the spatio-temporal structure of the fluctuations by studying the time-dependence of each $\mathbf{k}$-mode (although by doing so, we no longer have an ensemble of snapshots over which to average). In particular, we can expand the spatial Fourier components of any quantity $Q$ (see Section 2) as a Fourier series in time [^6] Because we are analyzing outputs over a finite temporal baseline $t_{\mathrm{sim}}$ (see Table 1), we have $\omega\equiv 2\pi n/t_{\mathrm{sim}}$ for $n\in\mathbb{Z}$. using
@@ 3.3 Two-Point Spatio-temporal Surface Density Statistics
In Figure 8 we illustrate the structure of this normalized spatio-temporal power as a function of both $k$ and $\omega$ for the linear fluctuations ($Q=\delta$) in the R8 simulation; analogous plots for the logarithmic fluctuations ($Q=s$) or the LGR4 simulations show similar features. Qualitatively, it is clear that modes with different $k$ have different temporal structure. However, across all $\mathbf{k}$-modes, the power peaks at $\omega=0$ and decreases monotonically with $\omega$ thereafter. This indicates that the time-dependence of an individual $\mathbf{k}$-mode’s amplitude is decorrelated from other modes with distinct $k$ — such correlation would lead to a peak at a nonzero characteristic frequency. Nonetheless, there are characteristic timescales in the problem: as we will illustrate below, these can be extracted from the width of the temporal power spectra.
@@
for a function $\psi_{Q}(\tilde{\omega})$ that encodes the temporal structure of each $\mathbf{k}$-mode (while $P_{Q}$ encodes the amplitude of the mode). Note that we have included an explicit extra factor of $\omega_{0}(k)t_{\mathrm{sim}}$ in the definition of $\psi_{Q}$ in comparison to the quantity plotted in Figure 8 so that the normalization of $\psi_{Q}$ is set by
@@
This guarantees that the total spatio-temporal power when integrated over both $\omega$ and $k$-space matches the total spatial power integrated over $k$-space. In the model specified by Equation 27 and Equation 28, the only free parameters associated with the temporal structure are therefore $\tau_{0}$ and $v_{\mathrm{eff}}$. We fit the values of these parameters from the simulation data by minimizing the variance between the values of the quantity $\psi_{Q}(\tilde{\omega})$ for different $k$ at each fixed $\tilde{\omega}$ value.
@@
Figure 9: The spatio-temporal spectra for the linear (left, panel (a)) and logarithmic (right, panel (b)) surface density fluctuations for the R8 (purple) and LGR4 (orange) simulations with the spatial spectrum measured in Figure 7 factored out to produce the quantity $\psi_{Q}$ (see Equation 28), which encapsulates the frequency structure. We plot $\psi_{Q}$ as a function of the normalized frequency $\tilde{\omega}\equiv\omega/\omega_{0}(k)$ (see Equation 27). The solid curves show an average of $\psi_{Q}(\tilde{\omega})$ over $k$ weighted by each $k$-value’s contribution to the power $P_{Q}(k)$, and the shaded region shows a ±$\pm 1\sigma$ range around that average. The dashed curve is a fit using the “squared-Lorentzian model” (see Equation 30).
In Figure 9, we show the resulting curves of $\psi_{Q}(\tilde{\omega})$ for the linear (panel (a)) and logarithmic (panel (b)) fluctuations in the R8 (purple) and LGR4 (orange) simulations: the solid curves are spatial-power-weighted averages over $k$, while the shaded region indicates a ±$\pm 1\sigma$ range around this weighted average. The relatively small extent of the ±$\pm 1\sigma$ region indicates that the differences in temporal structure among distinct $\mathbf{k}$-modes are well-captured by a “dispersion relation” of the form of Equation 27. Once again, despite the significant difference in galactic environment modeled by the R8 and LGR4 simulations, the two-point statistics presented in the Figure are remarkably similar. This suggests that the ISM’s response to feedback processes that shape the spatio-temporal structure are statistically consistent across different simulations, once normalized by the appropriate “dispersion relation” for the environment.
We find that the measured curves $\psi_{Q}(\tilde{\omega})$ presented in Figure 9 are well-fit by a “squared Lorentzian” function of the form
@@
we overplot this fit as a black dashed curve in the Figure. The model Equation 30 captures the plateau in power at the lowest $\tilde{\omega}$ values accurately, and falls within the ±$\pm 1\sigma$ interval around the $k$-averaged curve throughout the relevant $\tilde{\omega}$ range, although it slightly underpredicts the power at the highest $\tilde{\omega}$ in the R8 simulation and around $\tilde{\omega}\approx 1$ in the LGR4 simulation.
@@
where each $t_{j}$ is a “peak time” for the $j$-th structure at scale $k$ drawn uniformly from $[0,t_{\mathrm{sim}}]$. The Fourier transforms of individual terms in Equation 31 are proportional to $(\omega_{0}(k)^{2}+\omega^{2})^{-1}$, so that the total power from a set of such fluctuations would yield a squared Lorentzian for $\psi_{Q}$, as in the model of Equation 30. Thus, we may interpret the quantity $2\omega_{0}(k)^{-1}$ as a characteristic lifetime for a structure at scale $k$, which ranges from $2\tau_{0}\sim 10\,$Myr at the largest scales to $2/(kv_{\mathrm{eff}})\sim 0.3\,\mathrm{Myr}\times(k/(2\pi/10\,\mathrm{pc}))^{-1}$ at the smallest scales. Physically, these limits are simply the statement that at the smallest scales, the lifetime of a structure is of order the dynamical (i.e., turnover) timescale, while at the largest scales, structure lifetimes are determined by feedback processes.
Overall, compiling the results presented throughout Section 3, our model for the spatio-temporal power spectrum of the fluctuations takes the form
@@
×$\times\frac{4}{\omega_{0}(k)t_{\mathrm{sim}}}\left(1+\left(\frac{\omega}{\omega_{0}(k)}\right)^{2}\right)^{-2},$ (32)
@@
We emphasize that Section 3.3 provides a model for the amplitudes of individual fluctuation components. In order to realize a single snapshot (or a temporal sequence of snapshots), the relative phases of individual components must also be assigned. If the true phases were random, then since $s$ is well-described by a Gaussian pdf (see Figure 3), realizations of the surface density field could be generated using standard algorithms to produce Gaussian random fields (GRFs) by setting $Q=s$ in Section 3.3. However, as we shall show in Section 5.3, this procedure does not in fact succeed in recovering the true surface density, because the assumption of random phases is invalid for the nonlinear structure in the ISM.
@@
4 Volume Density and Gravitational Potential
@@
Although the midplane of the gravitational potential set by the dark matter halo and the stellar disk (Section 2) is fixed to $z=0$, panel (a) shows that the center of mass $z_{\mathrm{com}}(t)$ of the gas oscillates about this plane by several tens of pc in both simulations, as a result of asymmetry in feedback processes. Similarly, if there were only microphysical pressure supporting the disk against vertical gravity, in a steady-state the thickness would settle to a constant equilibrium value, but panel (b) demonstrates that $H$ varies by almost a factor of two in both simulations, due to the important role of large-scale, time-varying turbulence in supporting the disk (see E. C. Ostriker & C.-G. Kim, 2022, and references therein) [^8] Although both $z_{\mathrm{com}}(t)$ and $H$ have significant variations, simulations (at lower resolution than those shown here) show that a well-defined quasi-steady-state is maintained over nearly 1 Gyr timescales – see, e.g., Figure 1 of C.-G. Kim et al. 2023.. In both simulations, oscillations of $z_{\mathrm{com}}$ and $H$ occur on vertical gravitational oscillation periods of $\sim 80\,$Myr. For reference, we record the time-averaged value of the thickness and typical amplitude of the oscillations of both $z_{\mathrm{com}}$ and $H$ (in the form of their standard deviations as a function of time) in Table 4. In the R8 simulation, a sharp increase in $H$ and a decrease in $z_{\mathrm{com}}$ is apparent around $t\approx 300\,$Myr; this is due to a significant burst in feedback that occurs at that time.
@@
    Note that because the purpose of our models is to connect the fluctuations in the surface density field to those in the volume density and potential, we will restrict our approximations to the vertical quantities, and use the real, measured planar quantities (e.g., $\delta(\mathbf{x},t)$ and $\overline{\Sigma}(t)$) throughout.
@@
In Section 4.1 we begin by developing models for the vertical density and potential averaged over the in-plane variables ($k=0$), and in Section 4.2 we extend these models to characterize the potential fluctuations as a function of horizontal position $\mathbf{x}$ ($k>0$).
4.1 Mean Vertical Structure
@@
$\overline{\rho}_{\mathrm{m}}(z,t)=\frac{\overline{\Sigma}(t)}{2h_{\mathrm{m}}(t)}\zeta_{\mathrm{m}}(\tilde{z}),$ (43)
@@
Figure 11: The time average of the vertical profile $\zeta$ (defined in Equation 41) for the horizontally averaged ($k=0$) density in the R8 (purple, panel (a)) and LGR4 (orange, panel (b)) simulations. The shaded regions indicate ±$\pm 1\sigma$ error bars around the mean, which is indicated by the solid curve. In each panel, the fitted model $\zeta_{\mathrm{m}}$ constructed from a weighted combination of sech2 and exponential components is shown by a black dashed curve, with the individual sech2 and exponential components shown in grey dot-dashed and dotted curves respectively (see Equation 44).
In Figure 11 we plot the vertical profiles $\zeta\equiv 2h\overline{\rho}/\overline{\Sigma}$ as a function of $\tilde{z}$ measured in the R8 (purple) and LGR4 (orange) simulations, with the time averages shown as solid curves and the ±$\pm 1\sigma$ range about each average shown as a shaded region. For the R8 simulation, the relatively small deviations from the time-averaged profile indicates that there is indeed little vertical time-dependence outside of that encapsulated in $\tilde{z}$. In the LGR4 simulation, the stronger feedback produces more significant variability in moments of the vertical profile beyond $z_{\mathrm{com}}(t)$ and $H(t)$, resulting in greater snapshot-to-snapshot variation from the mean profile. The stronger feedback is also evident in the local maxima in the time-averaged curve at $|\tilde{z}|\approx 0.5$, which are produced when gas is consistently blown away from the midplane.
@@
The functional form of Equation 44 is physically well motivated. First, exponential density profiles are produced assuming hydrostatic equilibrium for an isothermal gas in a constant gravitational field, as we expect to be true at large $|\tilde{z}|$ (since only the hottest gas can extend far beyond the midplane, and the gravitational field of an infinite plane is independent of height). Second, $\mathrm{sech}^{2}$ density profiles are produced assuming hydrostatic equilibrium for an isothermal gas when the gas produces the confining gravitational field, as we expect to be true at small $|\tilde{z}|$ (L. Spitzer, 1942) [^9] The gas disk’s vertical structure also responds to the stellar disk’s vertical gravity, which can lead to a more Gaussian-like gas profile if the stellar disk is much thicker than that of the gas, but that is not the case for the current models.. Both the exponential and $\mathrm{sech}^{2}$ profiles are commonly applied in fitting observations (see, e.g., G. Jóhannesson et al. 2018), and are especially convenient because they allow for analytic solutions to the Poisson equation. In particular, the mean potential corresponding to the density profile specified by Equation 43 and Equation 44 is
@@
$+w\alpha_{\mathrm{e}}^{2}(e^{-|\tilde{z}|/\alpha_{\mathrm{e}}}-1)+w\alpha_{\mathrm{e}}|\tilde{z}|\bigg{]},$ (46)
@@
Table 5 summarizes the best-fit parameters $(w,\alpha_{\mathrm{s}},\alpha_{\mathrm{e}})$ for each simulation, and also provides the ratio of the effective thicknesses to the scale heights $h_{\mathrm{m}}/H\equiv\alpha_{w}$ for each model. The fits are overplotted as black dashed curves in Figure 11, with the contributions of the individual $\mathrm{sech}^{2}$ and exponential components in grey dot-dashed and dotted curves respectively. The top row of Figure 12 below compares the measured mean vertical potential $\overline{\phi}$ to the model $\overline{\phi}_{\mathrm{m}}$ produced using Section 4.1.
@@
R8 0.38$\pm$0.02 0.64$\pm$0.01 0.82$\pm$0.02 0.71$\pm$0.01
LGR4 0.07$\pm$0.06 1.02$\pm$0.02 2.6$\pm$1.8 1.1$\pm$0.16
Table 5: The model parameters for the fits to the $\zeta(\tilde{z})$ profiles shown in Figure 11, with columns as follows: (1) the simulation name, (2) the weight of the exponential term $w$, (3) the scale parameter of the $\mathrm{sech}^{2}$ term $\alpha_{\mathrm{s}}$, (4) the scale parameter of the exponential term $\alpha_{\mathrm{e}}$, and (5) the ratio of the model’s effective thickness to the mass-weighted RMS disk thickness, $\alpha_{w}=h/H$.
@@
4.2 Potential Fluctuation Vertical Structure
Expanding the gravitational potential and surface density fluctuations in $\mathbf{k}$ as in Section 2, from the Poisson equation (Equation 35), we find that the potential fluctuations for our model separable density given in Equation 38 have Fourier components

### v2

@@ III.2 Two-Point Spatial Surface Density Statistics
Although we have presented Figure 6 for just the $\delta$ field in the R8 simulation, we note that analogous plots of $P_{s}(\mathbf{k})$ for the R8 simulation, and for both $P_{\delta}(\mathbf{k})$ and $P_{s}(\mathbf{k})$ for the LGR4 simulations, look qualitatively the same. In particular, they all show a near-isotropic field with a slight preference for trailing waves. Even for the scales that deviate most strongly from isotropy, the variation in the power at a fixed $k$ but varied $(k_{x},k_{y})$ is well within $\sim 0.25\,$dex, and so for the remainder of this paper, for simplicity, we will approximate all fluctuation power spectra to be a function of the magnitude of the wavevector, $k$, alone (but see Section V.3 for further discussion).
Figure 7: Spatial power spectra of the linear fluctuations (left) and logarithmic fluctuations (right) for the R8 (purple) and LGR4 (orange) simulations. The solid curves of each color are the time-averaged spectra averaged over the azimuthal angle in $k$-space, while shaded regions show $\pm 1\sigma$ variations around each curve due to anisotropy (see, e.g., Figure 6) and time-dependence. The left and right shaded grey regions in each panel are excluded from fitting the power-law slope (indicated by the colored dashed curves) due to box-size and resolution effects respectively.
In Figure 7 we plot the power spectra in $\delta$ (left) and $s$ (right) for the R8 (purple) and LGR4 (orange) simulations as a function of $k$, calculated by averaging the spectra $P_{\delta}(\mathbf{k})$ and $P_{s}(\mathbf{k})$ over contours of constant $k$ (e.g., the white dashed circles in Figure 6). The solid curves all indicate the measured power spectra, while the shaded regions around them represent $\pm 1\sigma$ deviations in the values due to time-dependence in the fluctuation Fourier components and anisotropy. We see that the shaded regions are wider at smaller $k$, demonstrating the increased anisotropy present there, but the fractional variation remains small, illustrating that the approximation of isotropy is reasonable.
For both $P_{\delta}$ and $P_{s}$, the power spectra in both simulations are well-fit by power laws across a wide range of scales. We measure power law indices by fitting the slope of the curves over the range $kL\in[10,200]$; for smaller $kL$ box-size effects and anisotropy may become important, while for larger $kL$ there is contamination due to the proximity to the grid scale $\Delta x$ [^6] Incorporating smaller values of $kL$ in the fit does not substantially alter the results—the error bars on the slope are simply increased due to the larger fractional error in power measured for the largest scales, as indicated in the shaded region of Figure 7 and described above.. We demonstrate in Appendix A that our results for the largest scales, which carry the most power, are not affected by the simulation spatial resolution. Thus, our model for the power spectra takes the form
@@
We see from Figure 7 and Table 2 that despite the stark differences in the galactic environments and feedback strengths between the R8 and LGR4 simulations, their dimensionless fluctuation power spectra as functions of $kL$ have power law indices consistent with each other and even share similar normalizations. Because the power law spectral indices are sufficiently steep ($n_{\delta},n_{s}>2$), the power in these fluctuations is dominated by the largest scales. We speculate on the origin of the similarity across environments, as well as the broader implications of the steep power-law structure of the field in Section V.1.
III.3 Two-Point Spatio-temporal Surface Density Statistics
So far, we have focused on the spatial features of the fluctuations, and averaged the relevant statistics over time as a proxy for an ensemble average. However, we are also able to probe the spatio-temporal structure of the fluctuations by studying the time-dependence of each $\mathbf{k}$-mode (although by doing so, we no longer have an ensemble of snapshots over which to average). In particular, we can expand the spatial Fourier components of any quantity $Q$ (see Section II) as a Fourier series in time [^7] Because we are analyzing outputs over a finite temporal baseline $t_{\mathrm{sim}}$ (see Table 1), we have $\omega\equiv 2\pi n/t_{\mathrm{sim}}$ for $n\in\mathbb{Z}$. using
@@ III.3 Two-Point Spatio-temporal Surface Density Statistics
In Figure 8 we illustrate the structure of this normalized spatio-temporal power as a function of both $k$ and $\omega$ for the linear fluctuations ($Q=\delta$) in the R8 simulation; analogous plots for the logarithmic fluctuations ($Q=s$) or the LGR4 simulations show similar features. Qualitatively, it is clear that modes with different $k$ have different temporal structure. However, across all $\mathbf{k}$-modes, the power peaks at $\omega=0$ and decreases monotonically with $\omega$ thereafter. This indicates that the time-dependence of an individual $\mathbf{k}$-mode’s amplitude is decorrelated from other modes with distinct kk—such correlation would lead to a peak at a nonzero characteristic frequency. Nonetheless, there are characteristic timescales in the problem: as we will illustrate below, these can be extracted from the width of the temporal power spectra.
@@
[…]
@@
[…]
@@
[…]
In Figure 9, we show the resulting curves $\psi_{Q}(\tilde{\omega}) fo$r the linear (panel (a)) and logarithmic (panel (b)) fluctuations in the R8 (purple) and LGR4 (orange) simulations: the solid curves are spatial-power-weighted averages over $k$, while the shaded region indicates a $\pm 1\sigma$ range around this weighted average. The relatively small extent of the $\pm 1\sigma$ region indicates that the differences in temporal structure among distinct $\mathbf{k}$-modes are well-captured by a “dispersion relation” of the form of Equation 27. Once again, despite the significant difference in galactic environment modeled by the R8 and LGR4 simulations, the two-point statistics presented in the Figure are remarkably similar. This suggests that the ISM’s response to feedback processes that shape the spatio-temporal structure are statistically consistent across different simulations, once normalized by the appropriate “dispersion relation” for the environment.
[…]
@@
we overplot this fit as a black dashed curve in the Figure. The model Equation 30 captures the plateau in power at the lowest $\tilde{\omega}$ values accurately, and falls within the $\pm 1\sigma$ interval around the $k$-averaged curve throughout the relevant $\tilde{\omega}$ range, although it slightly underpredicts the power at the highest $\tilde{\omega}$ in the R8 simulation and around $\tilde{\omega}\approx 1$ in the LGR4 simulation.
@@
where each $t_{j}$ is a “peak time” for the $j$-th structure at scale $k$ drawn uniformly from $0,t_{\mathrm{sim}}]$. The Fourier transforms of individual terms in Equation 31 are proportional to $(\omega_{0}(k)^{2}+\omega^{2})^{-1}$, so that the total power from a set of such fluctuations would yield a squared Lorentzian for $\psi_{Q}$, as in the model of Equation 30. Thus, we may interpret the quantity $2\omega_{0}(k)^{-1}$ as a characteristic lifetime for a structure at scale $k$, which ranges from $2\tau_{0}\sim 10\,$Myr at the largest scales to $2/(kv_{\mathrm{eff}})\sim 0.3\,\mathrm{Myr}\times(k/(2\pi/10\,\mathrm{pc}))^{-1}$ at the smallest scales. Physically, these limits are simply the statement that at scales $\lambda\lesssim 2\pi v_{\mathrm{eff}}\tau_{0}\sim 300\,\mathrm{pc}$, the lifetime of a structure is of order the dynamical timescale associated with MHD waves and trans-sonic turbulence, while for larger scales, structure lifetimes are determined by feedback processes. We also note that these structure lifetimes are substantially shorter than the shear timescale $\sim(q\Omega)^{-1}\sim 30\,$Myr, so turbulence and feedback play a more significant role in dispersing structures than shear.
Overall, compiling the results presented throughout Section III, our model for the spatio-temporal power spectrum of the fluctuations takes the form
@@
$\times\frac{4}{\omega_{0}(k)t_{\mathrm{sim}}}\left(1+\left(\frac{\omega}{\omega_{0}(k)}\right)^{2}\right)^{-2},$ (32)
@@
We emphasize that Section III.3 provides a model for the amplitudes of individual fluctuation components. In order to realize a single snapshot (or a temporal sequence of snapshots), the relative phases of individual components must also be assigned. If the true phases were random, then since $s$ is well-described by a Gaussian pdf (see Figure 3), realizations of the surface density field could be generated e.g., by using standard algorithms to produce Gaussian random fields (GRFs) by setting $Q=s$ in Section III.3. We outline two methods for generating surface density fields in this way in Appendix B. However, as we shall show in Section V.3, this procedure does not in fact succeed in recovering the true surface density, because the assumption of random phases is invalid for the nonlinear structure in the ISM.
@@
IV Volume Density and Gravitational Potential
@@
Although the midplane of the gravitational potential set by the dark matter halo and the stellar disk (Section II) is fixed to $z=0$, panel (a) shows that the center of mass $z_{\mathrm{com}}(t)$ of the gas oscillates about this plane by several tens of pc in both simulations, as a result of asymmetry in feedback processes. Similarly, if there were only microphysical pressure supporting the disk against vertical gravity, in a steady-state the thickness would settle to a constant equilibrium value, but panel (b) demonstrates that $H$ varies by almost a factor of two in both simulations, due to the important role of large-scale, time-varying turbulence in supporting the disk (see E. C. Ostriker & C.-G. Kim, 2022, and references therein) [^9] Although both $z_{\mathrm{com}}(t)$ and $H$ have significant variations, simulations (at lower resolution than those shown here) show that a well-defined quasi-steady-state is maintained over nearly 1 Gyr timescales – see, e.g., Figure 1 of C.-G. Kim et al. 2023.. In both simulations, oscillations of $z_{\mathrm{com}}$ and $H$ occur on vertical gravitational oscillation periods of $\sim 80\,$Myr. For reference, we record the time-averaged value of the thickness and typical amplitude of the oscillations of both $z_{\mathrm{com}}$ and $H$ (in the form of their standard deviations as a function of time) in Table 4. In the R8 simulation, a sharp increase in $H$ and a decrease in $z_{\mathrm{com}}$ is apparent around $t\approx 300\,$Myr; this is due to a significant burst in feedback that occurs at that time.
@@
    That is, the model simply “inflates” all surface density structures to have the same vertical profile. Note that because the purpose of our models is to connect the fluctuations in the surface density field to those in the volume density and potential, we will restrict our approximations to the vertical quantities, and use the real, measured planar quantities (e.g., $\delta(\mathbf{x},t)$ and $\overline{\Sigma}(t)$) throughout.
@@
In Section IV.1 we begin by developing models for the vertical density and potential averaged over the in-plane variables ($k=0$), and in Section IV.2 we extend these models to characterize the potential fluctuations as a function of horizontal position $\mathbf{x}$ ($k>0$).
IV.1 Mean Vertical Structure
@@
[…]
@@
Figure 11: The time average of the vertical profile $\zeta$ (defined in Equation 41) for the horizontally averaged ($k=0$) density in the R8 (purple, panel (a)) and LGR4 (orange, panel (b)) simulations. The shaded regions indicate $\pm 1\sigma$ error bars around the mean, which is indicated by the solid curve. In each panel, the fitted model $\zeta_{\mathrm{m}}$ constructed from a weighted combination of sech2 and exponential components is shown by a black dashed curve, with the individual sech2 and exponential components shown in grey dot-dashed and dotted curves respectively (see Equation 44).
In Figure 11 we plot the vertical profiles $\zeta\equiv 2h\overline{\rho}/\overline{\Sigma}$ as a function of $\tilde{z}$ measured in the R8 (purple) and LGR4 (orange) simulations, with the time averages shown as solid curves and the $\pm 1\sigma$ range about each average shown as a shaded region. For the R8 simulation, the relatively small deviations from the time-averaged profile indicates that there is indeed little vertical time-dependence outside of that encapsulated in $\tilde{z}$. In the LGR4 simulation, the stronger feedback produces more significant variability in moments of the vertical profile beyond $z_{\mathrm{com}}(t)$ and $H(t)$, resulting in greater snapshot-to-snapshot variation from the mean profile. The stronger feedback is also evident in the local maxima in the time-averaged curve at $|\tilde{z}|\approx 0.5$, which are produced when gas is consistently blown away from the midplane.
@@
[…]
@@
[…]
@@
Table 5 summarizes the best-fit parameters $(w,\alpha_{\mathrm{s}},\alpha_{\mathrm{e}})$ for each simulation, and also provides the ratio of the effective thicknesses to the scale heights $h_{\mathrm{m}}/H\equiv\alpha_{w}$ for each model. The fits are overplotted as black dashed curves in Figure 11, with the contributions of the individual $\mathrm{sech}^{2}$ and exponential components in grey dot-dashed and dotted curves respectively. The top row of Figure 12 below compares the measured mean vertical potential $\overline{\phi}$ to the model $\overline{\phi}_{\mathrm{m}}$ produced using Section IV.1.
@@
[…]
@@
IV.2 Potential Fluctuation Vertical Structure
Expanding the gravitational potential and surface density fluctuations in $\mathbf{k}$ as in Section II, from the Poisson equation (Equation 35), we find that the potential fluctuations for our model separable density given in Equation 38 have Fourier components